Sowane’s Computer Training School in Sigatoka stocks workbooks with the



following characteristics:



Demand D = 18,500 units/year



Ordering cost S = $26/order



Holding cost H = $4/unit/year



a) Calculate the EOQ for the workbooks.



b) What are the annual holding costs for the workbooks?



c) What are the annual ordering costs?


1
Expert's answer
2021-11-10T14:53:04-0500

a) "EOQ = \sqrt{2*18,500*26/4} = 490.4" units.

b) The annual holding costs for the workbooks are:

490.4*4 = 1,961.63.

c) The annual ordering costs are:

490.4*26 = 12,750.61.